# Efficient quantum Gibbs samplers with Kubo–Martin–Schwinger detailed balance condition

### v1

@@ 1. Introduction
where $\left\lVert\cdot\right\rVert_{1}$ denotes the trace norm. It is worth mentioning that the quantum DBC is also instrumental for proving the finite mixing time (if it is the case), and its scaling with respect to $\beta$ and the system size [${}^{+}$10, KT13, ${}^{+}$23, RFA24]. However, due to the non-commutativity of operators in quantum mechanics, there is no unique definition of quantum DBC [${}^{+}$10, CM17, CM20].
@@ 1.2. Contribution
In parallel to Alicki’s characterization of Lindblad generators satisfying the GNS DBC, Fagnola, and Umanità [FU07, FU10] have prescribed the necessary and sufficient conditions for a quantum Markov semigroup to satisfy the KMS DBC. This leads to a set of conditions on the jump operators and the Hamiltonian for its corresponding Lindblad generator (also see [AC21]). Building upon these works, we introduce a family of quantum Gibbs samplers satisfying the KMS DBC. This includes the construction of [CKG23] as a special instance. In particular, $\gamma(\omega)$ can be chosen to be a discrete sum of $\delta$ functions, leading to a finite number of jump operators. In fact, it is sufficient to choose $\gamma(\omega)=\delta(\omega)$, which means a single jump operator if $\left\lvert\mathcal{A}\right\rvert=1$. Our jump operators can be constructed using the standard linear combination of unitaries (LCU) routine [${}^{+}$14, GSLW19]. As a result, our Lindblad dynamics can be efficiently simulated using any high-order simulation algorithms, including those in [LW23, DLL24]. In addition, we show that this new family of Gibbs samplers enables the selection of $f(t)$, whose Fourier transform is smooth and compactly supported. This approach simplifies the error analysis for controlling the discretization error, through the application of the Poisson summation formula. Table 1 compares the performance of a number of quantum Gibbs samplers based on the Lindblad dynamics.
@@ 1.3. Discussion and open questions
Although rigorous bounding of the mixing time has been achieved for certain quantum Gibbs samplers operating on commuting Hamiltonians [KB16, ${}^{+}$23], establishing the mixing time for non-commuting Hamiltonians at moderate or even low temperatures presents a substantial theoretical challenge. There are two interesting works along this line. The first is that Rouzé et al [RFA24] established the spectral gap of KMS detailed balanced Lindbladians for certain $k$-local Hamiltonians at high temperatures using Lieb-Robinson estimates. The analysis in [RFA24] may be applicable in our setting and we plan to investigate this in detail in a future work. On the other hand, Bakshi et al [BLMT24] demonstrated that at the high temperature, the Gibbs State of certain $k$-local Hamiltonian becomes a linear combination of tensor products of stabilizer states, can be prepared in polynomial time using randomized classical algorithms. This suggests that exploring the relationship between the complexity of Gibbs states and mixing times could be a fruitful avenue for future research.
@@ 2. Structures of detailed balanced Lindbladians
where Ψ()normal-$\Psi(\cdot)$ is completely positive with the Kraus representation:
@@ 2.2. KMS-detailed balanced generators
with the CP operator Ψ()normal-$\Psi(\cdot)$ admitting the Kraus representation (2.2) and the operator
@@
[…]
@@ 3.1. Quantum Gibbs samplers via KMS-detailed balanced Lindbladian
To ensure that the constructed Lindblad dynamics (3.8) eventually relaxes to the desired Gibbs state, by 3, we should carefully choose the coupling operators $A^{a}$ and weighting functions $q^{a}$ such that the resulting $\{L^{a}\}$ and $G$ satisfy Eq. 2.5. This is always possible, due to the finite dimensionality of the system. Moreover, it is known [${}^{+}$10] that for a primitive KMS detailed balanced QMS, the mixing time (1.3) can be characterized by the spectral gap of the Lindbladian. The recent work [RFA24] estimated the spectral gap of the efficient quantum Gibbs sampler in [CKG23] in the high-temperature regime, by mapping the Lindbladian $\mathcal{L}_{\beta}$ with KMS DBC to a Hamiltonian ~$\widetilde{\mathcal{L}}_{\beta}:=\sigma_{\beta}^{-1/4}\mathcal{L}_{\beta}(\sigma_{\beta}^{1/4}X\sigma_{\beta}^{1/4})\sigma_{\beta}^{-1/4}$ and then analyzing its spectral properties with perturbation theory. The extension of such a mixing time analysis framework to our case with the optimal selection of $\{(A^{a},q^{a})\}_{a\in\mathcal{A}}$ will be discussed in a forthcoming work.
@@ 3.2. Choice of the weighting function $q(\nu)$
(3.14) $(M-1)\tau=\Omega\left((\beta A_{u}+A_{w})\log\left(\beta A_{u}+A_{w}\right)^{s}\right)\,,$
@@
Here Ω()normal-$\Omega(\cdot)$, $\mathcal{O}(\cdot)$ absorbs some constant depending on $s$.
@@ Appendix D Quadrature error analysis
Next, by Eq. D.1, a direct computation via change of variable $t^{s-1}=u$ gives
(D.11) $\int_{T}^{\infty}\left|f(t)\right|\,\mathrm{d}t=\mathcal{O}\left(A^{2}_{q}S\int_{T}^{\infty}u^{s-1}\exp\left(-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}u\right)\,\mathrm{d}u\right)\,,\quad T>0\,.$
@@
(D.12) $T_{f}(A_{u},A_{w}):=\inf\left\{T>0\,;\ u^{s-1}\leq\exp\left(\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}u\right)\ \text{for any $u\geq T$}\right\},$
@@
$T_{f}(A_{u},A_{w})/\log(T_{f}(A_{u},A_{w}))=\Theta\left((\beta A_{u}+A_{w})^{1/s}\right)\,.$
@@
(D.13) $T_{f}(A_{u},A_{w})=\Theta\left((\beta A_{u}+A_{w})^{1/s}\log\left(\beta A_{u}+A_{w}\right)\right)\,.$
@@
(D.14) $\int_{\mathbb{R}}\left|f(t)\right|\,\mathrm{d}t$ $=\int_{-T_{f}}^{T_{f}}|f(t)|\,\mathrm{d}t+\mathcal{O}\left(A^{2}_{q}S\int_{T_{f}}^{\infty}\exp\left(-\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}u\right)\,\mathrm{d}u\right)$
$=\mathcal{O}\left((A_{q}C_{1,u}+A^{2}_{q}SA_{w})\log T_{f}+A^{2}_{q}S\underbrace{(\beta A_{u}+A_{w})^{1/s}\exp\left(-\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}T_{f}\right)}_{(\beta A_{u}+A_{w})^{1/s-\Theta(s)}\leq 1}\right)$
$=\mathcal{O}\left((A_{q}C_{1,u}+A^{2}_{q}SA_{w})\log(\beta A_{u}+A_{w})\right)\,.$
This gives Eq. D.3.
Estimate the integral of $g$. Similarly, by Eq. D.6 with 15 and $|\tanh((\nu^{\prime}-\nu)/4)|\leq 1$, we have
$\left\lVert g(t,t^{\prime})\right\rVert_{L^{\infty}(\mathbb{R}^{2})}\leq\frac{1}{2\pi^{2}}A_{q}^{4}S^{2}\,.$
A direct computation gives, for a fixed $\nu^{\prime}\in\mathbb{R}$,
@@
(D.16) $T_{g}(A_{u},A_{w}):=\inf\left\{T>0\,;\ u^{2s-1}\leq\exp\left(\frac{s}{2(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}u\right)\ \text{for any $u\geq T$}\right\},$
@@
$T_{g}(A_{u},A_{w})=\Theta\left((2\beta A_{u}+2A_{w}+\beta)^{1/s}\log\left(2\beta A_{u}+2A_{w}+\beta\right)\right)\,.$
@@
$=\int_{-T_{g}}^{T_{g}}\int_{-T_{g}}^{T_{g}}|g(t,t^{\prime})|\,\mathrm{d}t\,\mathrm{d}t^{\prime}+\mathcal{O}\left(A^{4}_{q}S^{2}\int_{T_{g}}^{\infty}\exp\left(-\frac{s}{(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}u\right)u^{2s-1}\,\mathrm{d}u\right)$
@@
Let $T_{f}(A_{u},A_{w})$ be the constant defined as in Eq. D.12. Then, when $(M-1)\tau\geq T_{f}(A_{u},A_{w})^{s}$, we can compute, by (D.18),
@@
Estimate for the coherent term $G$. Again by 30 with similar estimates as in (D.19) for $L_{a}$, we obtain, when $(M-1)\tau>T_{g}(A_{u},A_{w})^{s}$,

### v2

@@ 1. Introduction
where $\left\lVert\cdot\right\rVert_{1}$ denotes the trace norm. It is worth mentioning that the quantum DBC is also instrumental for proving the finite mixing time (if it is the case), and its scaling with respect to $\beta$ and the system size [TKR+10, KT13, BCG+23, RFA24]. However, due to the non-commutativity of operators in quantum mechanics, there is no unique definition of quantum DBC [TKR+10, CM17, CM20].
@@ 1.2. Contribution
In parallel to Alicki’s characterization of Lindblad generators satisfying the GNS DBC, Fagnola, and Umanità [FU07, FU10] have prescribed the necessary and sufficient conditions for a quantum Markov semigroup to satisfy the KMS DBC. This leads to a set of conditions on the jump operators and the Hamiltonian for its corresponding Lindblad generator (also see [AC21]). Building upon these works, we introduce a family of quantum Gibbs samplers satisfying the KMS DBC. This includes the construction of [CKG23] as a special instance. In particular, $\gamma(\omega)$ can be chosen to be a discrete sum of $\delta$ functions, leading to a finite number of jump operators. In fact, it is sufficient to choose $\gamma(\omega)=\delta(\omega)$, which means a single jump operator if $\left\lvert\mathcal{A}\right\rvert=1$. Our jump operators can be constructed using the standard linear combination of unitaries (LCU) routine [BCC+14, GSLW19]. As a result, our Lindblad dynamics can be efficiently simulated using any high-order simulation algorithms, including those in [LW23, DLL24]. In addition, we show that this new family of Gibbs samplers enables the selection of $f(t)$, whose Fourier transform is smooth and compactly supported. This approach simplifies the error analysis for controlling the discretization error, through the application of the Poisson summation formula. Table 1 compares the performance of a number of quantum Gibbs samplers based on the Lindblad dynamics.
@@ 1.3. Discussion and open questions
Although rigorous bounding of the mixing time has been achieved for certain quantum Gibbs samplers operating on commuting Hamiltonians [KB16, BCG+23], establishing the mixing time for non-commuting Hamiltonians at moderate or even low temperatures presents a substantial theoretical challenge. There are two interesting works along this line. The first is that Rouzé et al [RFA24] established the spectral gap of KMS detailed balanced Lindbladians for certain $k$-local Hamiltonians at high temperatures using Lieb-Robinson estimates. The analysis in [RFA24] may be applicable in our setting and we plan to investigate this in detail in a future work. On the other hand, Bakshi et al [BLMT24] demonstrated that at the high temperature, the Gibbs State of certain $k$-local Hamiltonian becomes a linear combination of tensor products of stabilizer states, can be prepared in polynomial time using randomized classical algorithms. This suggests that exploring the relationship between the complexity of Gibbs states and mixing times could be a fruitful avenue for future research.
@@ 2. Structures of detailed balanced Lindbladians
where $\Psi(\cdot)$ is completely positive with the Kraus representation:
@@ 2.2. KMS-detailed balanced generators
with the CP operator $\Psi(\cdot)$ admitting the Kraus representation (2.2) and the operator
@@
(2.34) $\mathcal{L}(X)=\sum_{j\in\mathcal{J}}\mathcal{L}_{j}(X)\quad\text{with}\quad\mathcal{L}_{j}(X)=2e^{-\omega_{j}/2}\left(L_{j}^{\dagger}XL_{j}-\frac{1}{2}\left\{L_{j}^{\dagger}L_{j},X\right\}\right)\,,$
@@ 3.1. Quantum Gibbs samplers via KMS-detailed balanced Lindbladian
To ensure that the constructed Lindblad dynamics (3.8) eventually relaxes to the desired Gibbs state, by 3, we should carefully choose the coupling operators $A^{a}$ and weighting functions $q^{a}$ such that the resulting $\{L^{a}\}$ and $G$ satisfy Eq. 2.5. This is always possible, due to the finite dimensionality of the system. Moreover, it is known [TKR+10] that for a primitive KMS detailed balanced QMS, the mixing time (1.3) can be characterized by the spectral gap of the Lindbladian. The recent work [RFA24] estimated the spectral gap of the efficient quantum Gibbs sampler in [CKG23] in the high-temperature regime, by mapping the Lindbladian $\mathcal{L}_{\beta}$ with KMS DBC to a Hamiltonian ~$\widetilde{\mathcal{L}}_{\beta}:=\sigma_{\beta}^{-1/4}\mathcal{L}_{\beta}(\sigma_{\beta}^{1/4}X\sigma_{\beta}^{1/4})\sigma_{\beta}^{-1/4}$ and then analyzing its spectral properties with perturbation theory. The extension of such a mixing time analysis framework to our case with the optimal selection of $\{(A^{a},q^{a})\}_{a\in\mathcal{A}}$ will be discussed in a forthcoming work.
@@ 3.2. Choice of the weighting function $q(\nu)$
(3.14) $(M-1)\tau=\Omega\left((\beta A_{u}+A_{w})\log\left(\beta A_{u}+A_{w}\right)\right)\,,$
@@
Here $\Omega(\cdot)$, $\mathcal{O}(\cdot)$ absorbs some constant depending on $s$.
@@ 4. Recovery of the Gibbs sampler in [CKG23]
Conflict of interest statment
Authors have no conflict of interest to declare.
@@ Appendix D Quadrature error analysis
Next, by Eq. D.1, a direct computation via change of variable $t^{1/s}=u$ gives
(D.11) $\int_{T}^{\infty}\left|f(t)\right|\,\mathrm{d}t=\mathcal{O}\left(A^{2}_{q}S\int_{T^{1/s}}^{\infty}u^{s-1}\exp\left(-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}u\right)\,\mathrm{d}u\right)\,,\quad T>0\,.$
@@
(D.12) $T_{f}(A_{u},A_{w}):=\inf\left\{T>0\,;\ u^{s-1}\leq\exp\left(\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}u\right)\ \text{for any $u\geq T^{1/s}$}\right\},$
@@
$T_{f}(A_{u},A_{w})/\log(T_{f}(A_{u},A_{w}))=\Theta\left(\beta A_{u}+A_{w}\right)\,.$
@@
(D.13) $T_{f}(A_{u},A_{w})=\Theta\left((\beta A_{u}+A_{w})\log\left(\beta A_{u}+A_{w}\right)\right)\,.$
@@
(D.14) $\int_{\mathbb{R}}\left|f(t)\right|\,\mathrm{d}t$ $=\int_{-T^{k}_{f}}^{T^{k}_{f}}|f(t)|\,\mathrm{d}t+\mathcal{O}\left(A^{2}_{q}S\int_{T_{f}^{k/s}}^{\infty}\exp\left(-\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}u\right)\,\mathrm{d}u\right)$
$=\mathcal{O}\left((A_{q}C_{1,u}+A^{2}_{q}SA_{w})\log T_{f}+A^{2}_{q}S\underbrace{(\beta A_{u}+A_{w})^{1/s}\exp\left(-\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}T_{f}^{k/s}\right)}_{\text{$o(1)$ as $\beta A_{u}+A_{w}\to\infty$}}\right)$
$=\mathcal{O}\left((A_{q}C_{1,u}+A^{2}_{q}SA_{w})\log(\beta A_{u}+A_{w})\right),$
for any positive $k>1$. This gives Eq. D.3.
Estimate the integral of $g$. Similarly, by Eq. D.6 with 15 and $|\tanh((\nu^{\prime}-\nu)/4)|\leq 1$, we have $\left\lVert g(t,t^{\prime})\right\rVert_{L^{\infty}(\mathbb{R}^{2})}\leq\frac{1}{2\pi^{2}}A_{q}^{4}S^{2}$. A direct computation gives, for a fixed $\nu^{\prime}\in\mathbb{R}$,
@@
(D.16) $T_{g}(A_{u},A_{w}):=\inf\left\{T>0\,;\ u^{2s-1}\leq\exp\left(\frac{s}{2(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}u\right)\ \text{for any $u\geq T^{1/s}$}\right\},$
@@
$T_{g}(A_{u},A_{w})=\Theta\left((2\beta A_{u}+2A_{w}+\beta)\log\left(2\beta A_{u}+2A_{w}+\beta\right)\right)\,.$
@@
$=\int_{-T_{g}}^{T_{g}}\int_{-T_{g}}^{T_{g}}|g(t,t^{\prime})|\,\mathrm{d}t\,\mathrm{d}t^{\prime}+\mathcal{O}\left(A^{4}_{q}S^{2}\int_{T_{g}^{1/s}}^{\infty}\exp\left(-\frac{s}{(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}u\right)u^{2s-1}\,\mathrm{d}u\right)$
@@
Let $T_{f}(A_{u},A_{w})$ be the constant defined as in Eq. D.12. Then, when $(M-1)\tau\geq T_{f}(A_{u},A_{w})$, we can compute, by (D.18),
@@
Estimate for the coherent term $G$. Again by 30 with similar estimates as in (D.19) for $L_{a}$, we obtain, when $(M-1)\tau>T_{g}(A_{u},A_{w})$,